OpenStudy (anonymous):

write 2 seperate equations. round to the nearest hundredth. -5+-2 sqrt3 over 3

OpenStudy (anonymous):

Is the 3 dividing all terms?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[\frac{-5\pm2\sqrt{3}}{3}\] First equation:\[\frac{-5+2\sqrt{3}}{3}\] Second equation:\[\frac{-5-2\sqrt{3}}{3}\]

OpenStudy (anonymous):

Enter this into your calculator to find the answers and round.

OpenStudy (anonymous):

what if you have to show your work

OpenStudy (anonymous):

Well you used the quadratic formula so show working for that but to get the answers rounded to the nearest hundredth you need to just type it into a calculator.

OpenStudy (anonymous):

Show how you separated the quadratic like I did above and then work out the answer :)

OpenStudy (anonymous):